# Multi-period Stochastic Network Design for Combined Natural Gas and Hydrogen Distribution

### v3

@@ Abstract
Hydrogen is produced from water using renewable electricity. Unlike electricity, hydrogen can be stored in large quantities for long periods. This storage ability acts as a green battery, allowing solar and wind energy to be generated and used at different times. As a result, green hydrogen plays a central role in facilitating a climate-neutral economy. However, the logistics for hydrogen are complex. As new hydrogen pipelines are developed for hydrogen, there is a trend toward repurposing the natural gas network for hydrogen, due to its economic and environmental benefits. Yet, a rapid conversion could disrupt the balance of natural gas supply and demand. Furthermore, technical and economic developments surrounding the transition contribute additional complexity, which introduces uncertainty in future supply and demand levels for both commodities. To address these challenges, we introduce a multi-period stochastic network design problem for the transition of a natural gas pipeline network into a green hydrogen pipeline network. We develop a progressive-hedging-based metaheuristic to solve the problem. Results demonstrate our matheuristic is efficient, both in computation time and in solution quality. We show that factoring in uncertainty avoids premature expansion and ensures the development of an adequate pipeline network meeting long-term needs. In a case study in the Northern Netherlands for Hydrogen Energy Applications in Valley Environments for Northern Netherlands initiative, we focus on two key scenarios: local production and importation, exploring their impacts on performance indicators. Our case insights exemplify the solid foundation for strategic decision-making in energy transitions through our approach.
@@ 1 Introduction
To solve this problem, we design an efficient matheuristic based on progressive hedging. We augment this matheuristic with several acceleration techniques. On well-calibrated benchmark instances, we show that our matheuristic outperforms a compact mixed-integer programming (MIP) formulation that is solved by Gurobi. Using the matheuristic, we study the structure of our optimal solutions. Among others, considering the stochastic nature of supply and demand leads to (i) an increased deployment of hydrogen infrastructure at the end of the planning horizon, reflecting a strategic investment that enhances the network’s capacity and efficiency, (ii) a later adoption of hydrogen infrastructure, thereby reducing the risk of overcapacity and underutilization through better-informed planning, and (iii), a higher need for installation capacity in the peak adaptation periods as more pipelines need to be built and retrofitted in a shorter amount of time.
@@ 3.1 System Description
We consider an energy transition network design problem on a finite planning horizon with periods $t\in\mathscr{T}\coloneqq\{0,1,\ldots,T-1\}$. We consider a directed network $\mathscr{G}=(\mathscr{N},\mathscr{A})$, where $\mathscr{N}$ is the node set and $\mathscr{A}$ the arc set.
At period $t=0$, we assume the network is constructed so that it enables matching demand and supply in the existing energy market. For this, we consider multiple commodities $k\in\mathscr{K}\coloneqq\{1,2,\ldots,K\}$ that model the different energy types. Each node $i\in\mathscr{N}\coloneqq\{1,2,\ldots,N\}$ has a net supply quantity for each commodity, which is negative for demand nodes and zero for intermediate nodes. For some commodities, there may be no market initially, which means that their net supply quantities are zero for each node. This also means that an intermediate node for a commodity could become a supply or a demand node in later periods of the time horizon, for example when an electrolysis facility for hydrogen is built at that node. We assume each of these nodes is included in the initial set of nodes $\mathscr{N}$, and thus the node set does not change over periods of the horizon. One node may be shared by multiple commodities, for example, a supplier node for one commodity may be a demand node for another commodity.
@@
We model this problem as a two-stage mixed integer with continuous recourse. This is in line with the observed complexity of stochastic network design models in the extant literature. Indeed, researchers often employ a two-stage approximation of the underlying multi-stage structure (Wang et al., 2019; Jiang et al., 2021). This simplification is also observed in other stochastic literature, such as scenario tree generation (Heitsch and Römisch, 2009), and rolling-horizon-based methods (Cavagnini et al., 2022), both of which involve approximations of multi-stages with two-stage models. In line with these approaches, our two-stage model implements a finite set of scenarios at the second stage.
In the first stage, we decide on which arcs to construct for the entire planning horizon, and simultaneously assign a specific commodity to each newly constructed arc, as well as decide on converting existing arcs from one commodity to another. We assume there is sufficient lead time between model implementation and the planning horizon to ensure the feasibility of strategic pipeline planning. This allows for all necessary preparatory activities to be completed, ensuring that pipeline construction can proceed as scheduled within the model’s planning horizon.
In the second stage, we model the uncertainty using a finite set of scenarios. We optimize for each scenario the distribution of each commodity to ensure supply and demand are balanced. We further incorporate in our model the flexibility to handle unmet demand by penalizing it, simulating alternative transport options such as trucks. To address practical questions regarding seasonal storage, our model adopts a two-season framework per year —winter and summer–— as commonly practiced in the HEAVENN region, where, for example, excess supply is allowed to be stored and carried over to the next seasons. The details of this structure, as well as how we generate our instance parameters in our study are given in A. The goal is to find a network design that minimizes the construction cost and the expected distribution cost.
@@ 3.2 Mixed Integer Programming Formulation
We model uncertainty with finitely many scenarios $\omega\in\Omega\coloneqq\{1,2,\ldots,W\}$. Each scenario $\omega$ has a probability of occurrence of $p_{\omega}$. Under scenario $\omega$, $s_{it}^{k\omega}$ represents the net supply of node $i$ for commodity $k$ in period $t$. Let $x_{at}^{k\omega}$ be the flow amount of commodity $k$ on arc $a$ on period $t$ under scenario $\omega$, which costs $c_{at}^{k}$ per unit, which we set linearly proportional to the arc length. Let $I_{it}^{k\omega}$ be the inventory level of node $i$ for commodity $k$ at the beginning of period $t\in\mathscr{T}\cup\{T\}$ under scenario $\omega$, which cannot exceed the inventory capacity of $C_{it}^{k\omega}$. Holding costs, which typically include depreciation, risk of damage, and opportunity costs (Fallah, 2011), are less accounted for in hydrogen applications (HEAVENN, 2022). Thus, we assume no inventory holding costs in this study. However, the model can be adjusted to include these costs if necessary, without affecting the methodology presented in this paper. Let $e_{it}^{k\omega}$ be the amount of unsatisfied net demand in node $i$ for commodity $k$ in period $t$ under scenario $\omega$, which is penalized with a cost of $h$ per unit. This penalty may, for example, represent the variable cost of emergency transport by truck. Let the sets $\delta^{+}(i)$ and $\delta^{-}(i)$ represent the arc sets that enter and leave node $i$, respectively.
@@
The objective function minimizes the total cost of arc construction, arc conversion, the expected cost of arc flow, and the expected penalty cost. Constraint (2) ensures the inventory levels are balanced between periods according to the net supply, as well as the inflow and outflow of the commodity. Constraint (3) ensures that the inventory respects the capacity. Constraint (4) ensures that the flow capacity of an arc is not exceeded, and is only available for the assigned commodity. Constraint (5) ensures each arc can be built once in the planning horizon. Constraint (6) ensures that at the period an arc is constructed, it is assigned to its corresponding commodity. In the following periods, the arc may be converted for the use of another commodity. For those periods, Constraint (7) satisfies that each constructed arc is assigned to one of the commodities. We link the variable $z$ with the conversion variable $b$ by Constraint (8), ensuring the conversion cost is reflected in the objective function value. Each arc is restricted to be converted at most once during the planning horizon by Constraint (9). The initial networks for each commodity are created by Constraint (10). Finally, with Constraint (11) we make sure that the initial inventories are assigned to the start inventory values of $o_{i}^{k}$. The remaining constraints model the decision variables’ domains.
@@ 4 Progressive Hedging Based Matheuristic
$b_{at}^{k}\in\{0,1\}$ $\forall a\in\mathscr{A},k\in\mathscr{K},\omega\in\Omega,t\in\mathscr{T}.$ (40)
As proposed by Rockafellar and Wets (1991), we relax the non-anticipativity constraints by removing Constraints (35) - (40), allowing the model for each $\omega\in\Omega$ to be solved individually. The objective for each scenario $\omega$ is, then, modified with the following Lagrangian coefficients:
@@
Here, $\lambda$, $\delta$, and $\pi$ are Lagrangian multipliers for decision variables $y$, $b$, and $z$ respectively. The ratio $\rho$ is used to enforce non-anticipativity by penalizing the deviations of decision variables from their average values across scenarios, as well as acting as a step size for the updates on multipliers. The average values of decision variables over scenarios, $\overline{y}$, $\overline{b}$, and $\overline{z}$, are calculated from the previous iteration of the algorithm. Since all the first-stage variables are binary in our model, the quadratic objective function (41) would be equal to the following linearized objective function for all binary solutions:
@@
Algorithm 1 Progressive Hedging.
@@
We apply four accelerating techniques to the base PHA in our solution method. The first acceleration technique is to consider bundles of scenarios instead of single-scenario subproblems. This results in solving two-stage recourse models (i.e., subject to the scenarios in the bundle) in the PHA, instead of solving deterministic single scenario problems. It has been shown that for well-calibrated bundle sizes, this reduces computation time without affecting the quality too much (Crainic et al., 2014). We use a set of bundles $\Xi$ where each scenario $\omega\in\Omega$ is in exactly one bundle $\xi\in\Xi$. The probability of a bundle is assigned as $p_{\xi}=\sum_{\omega\in\xi}p_{\omega}$. The objective function of the two-stage recourse model for bundle $\xi$ is given below in Equation (43).
@@
We use a similar acceleration technique with two main differences. Firstly, we enable fixing variables even for the cases where they are not equal for all scenarios but for the majority of scenarios. If the total probability of scenarios where the decision variable shares a common value exceeds an input probability parameter $p_{\textsc{h}}$, we fix the corresponding decision variables. Secondly, we observe that even if some decision variables do not share the same value, we may see similarities between the values of decision variables in terms of our problem context. Thus, we implement a problem-specific convergence rule that does not depend on individual decision variables as in Watson and Woodruff (2011). Instead, the convergence condition depends on subsets of decision variables. For example, consider a case where one arc is built for the majority of scenarios from node $i_{1}$ to $i_{2}$ for commodity $k$. However, in each scenario, a different arc with a different capacity may be built, or the same arc may be built in different time periods. In this case, the individual values of $y_{at}^{k\omega}$ suggest the algorithm is far from being converged, but the following would be satisfied for the majority of scenarios:
@@
To implement this logic of convergence, we create the acceleration technique given in Algorithm 2. This algorithm checks for Constraint (44) which is defined for decision variable $y$. We observe in preliminary experiments that the complexity of our model is mainly driven by the $y$ variables. As such, there is no need to consider the same logic for the variables $z$ and $b$. Indeed, they can be found rather easily for a fixed set of $y$ variables. This procedure is run once at each iteration of PHA.
Algorithm 2 addHeuristic
@@
The third acceleration technique enables early termination of the PHA. As mentioned above, the majority of first-stage decision variables converge rather fast in the PHA. Often, the algorithm spends the majority of the iterations for some small discrepancies in the last decision variables. In order to accelerate, we terminate the algorithm before it is fully converged, and solve the rest of the problem with a MIP solver. One way of doing this would be fixing a subset of first-stage decision variables, and then solving the MIP for the rest. However, we implement a similar idea to the heuristic constraints mentioned above and create heuristic constraints without fixing individual decision variables. It has the same heuristic constraint structure of Algorithm 2, but for early convergence, we also check the cases where Constraint (44) sums to zero. At each iteration, Algorithm 3 is executed in our method. Once the number of these heuristic constraints reaches the target stopping condition, we solve a MIP for the base model including these constraints.

### v4

@@ Abstract
Hydrogen is produced from water using renewable electricity. Unlike electricity, hydrogen can be stored in large quantities for long periods. This storage ability acts as a green battery, allowing solar and wind energy to be generated and used at different times. As a result, green hydrogen plays a central role in facilitating a climate-neutral economy. However, the logistics for hydrogen are complex. As new pipelines are developed for hydrogen, there is a trend toward repurposing the natural gas network for hydrogen, due to its economic and environmental benefits. Yet, a rapid conversion could disrupt the balance of natural gas supply and demand. Furthermore, technical and economic developments surrounding the transition contribute additional complexity, which introduces uncertainty in future supply and demand levels for both commodities. To address these challenges, we introduce a multi-period stochastic network design problem for the transition of a natural gas pipeline network into a green hydrogen pipeline network. We develop a progressive-hedging-based metaheuristic to solve the problem. Results demonstrate our matheuristic is efficient, both in computation time and in solution quality. We show that factoring in uncertainty avoids premature expansion and ensures the development of an adequate pipeline network meeting long-term needs. In a case study in the Northern Netherlands for Hydrogen Energy Applications in Valley Environments for Northern Netherlands initiative, we focus on two key scenarios: local production and importation, exploring their impacts on performance indicators. Our case insights exemplify the solid foundation for strategic decision-making in energy transitions through our approach.
@@ 1 Introduction
To solve this problem, we design an efficient matheuristic based on progressive hedging. We augment this matheuristic with several acceleration techniques. Exact methods like Benders decomposition are also common for stochastic network design, but they face significant computational challenges with the multi-period complexity and scale of our problem, and are therefore not pursued in this study. On well-calibrated benchmark instances, we show that our matheuristic outperforms a compact mixed-integer programming (MIP) formulation that is solved by Gurobi. Using the matheuristic, we study the structure of our optimal solutions. Among others, considering the stochastic nature of supply and demand leads to (i) an increased deployment of hydrogen infrastructure at the end of the planning horizon, reflecting a strategic investment that enhances the network’s capacity and efficiency, (ii) a later adoption of hydrogen infrastructure, thereby reducing the risk of overcapacity and underutilization through better-informed planning, and (iii), a higher need for installation capacity in the peak adaptation periods as more pipelines need to be built and retrofitted in a shorter amount of time.
@@ 3.1 System Description
We consider an energy transition network design problem on a finite planning horizon with a directed network $\mathscr{G}=(\mathscr{N},\mathscr{A})$, where $\mathscr{N}$ is the node set and $\mathscr{A}$ the arc set. We consider a setting with periods $t\in\mathscr{T}\coloneqq\{0,1,\ldots,T-1\}$. At each time period $t$, decisions regarding infrastructure (construction or conversion) can be implemented and network flows are to be determined. For the purpose of this study, we define each period as a six-month interval (e.g., a winter or summer season), a discretization chosen to capture seasonal variations in energy supply and demand. However, alternative time discretizations can be adopted without altering the model.
At period $t=0$, we assume the network is constructed so that it enables matching demand and supply in the existing energy market. For this, we consider multiple commodities $k\in\mathscr{K}\coloneqq\{1,2,\ldots,K\}$ that model the different energy types, such as natural gas or hydrogen. Each node $i\in\mathscr{N}\coloneqq\{1,2,\ldots,N\}$ has a net supply quantity for each commodity, which is negative for demand nodes and zero for intermediate nodes. The demand may be satisfied by any supplier within the same commodity network, with no fixed origin-destination pair as seen in classical multi-commodity network design. This assumption reflects a national-level energy system without contract-based routing. If needed, the model can accommodate more detailed commodity definitions, such as separate purity grades of hydrogen for distinct use in industrial or transportation applications, each represented as a distinct commodity. For some commodities, there may be no market initially, which means that their net supply quantities are zero for each node. This also means that an intermediate node for a commodity could become a supply or a demand node in later periods of the time horizon, for example when an electrolysis facility for hydrogen is built at that node. We assume each of these nodes is included in the initial set $\mathscr{N}$, and thus the node set does not change over periods of the horizon. One node may be shared by multiple commodities, for example, a supplier node for one commodity may be a demand node for another commodity.
@@
We model this problem as a two-stage mixed-integer program with continuous recourse. This is consistent with the complexity observed in stochastic network design problems, where two-stage approximations are commonly used to replace multi-stage formulations (Wang et al., 2019; Jiang et al., 2021). Similar simplifications are found in scenario tree generation (Heitsch and Römisch, 2009) and rolling-horizon methods (Cavagnini et al., 2022), which often emulate multi-stage dynamics through tractable two-stage constructs. More importantly, high-pressure transmission pipelines, whether newly constructed or repurposed, typically require 24–60 months from final investment decision to commissioning (European Network of Transmission System Operators for Gas (2024), ENTSOG). Such long lead times limit the practicality of multi-stage recourses and thus favor a two-stage representation in which infrastructure commitments are fixed prior to the planning horizon (Basciftci et al., 2019). We assume that the model is implemented with sufficient advance notice to allow all preparatory activities to be completed, ensuring that construction can proceed as scheduled within the planning horizon. If needed, the two-stage model can always be run in a rolling-horizon fashion, but given the long lead time between planning and execution, we limit our scope to the canonical two-stage mixed-integer program with continuous recourse.
In the first stage, we decide on which arcs to construct for the entire planning horizon, and simultaneously assign a specific commodity to each newly constructed arc, as well as decide on converting existing arcs from one commodity to another. In the second stage, we model the uncertainty using a finite set of scenarios. We optimize for each scenario the distribution of each commodity to ensure supply and demand are balanced. We further incorporate in our model the flexibility to handle unmet demand by penalizing it, simulating alternative transport options such as trucks. To address practical questions regarding seasonal storage, our model adopts a two-season framework per year —winter and summer— as commonly practiced in the HEAVENN region, where, for example, excess supply is allowed to be stored and carried over to the next seasons. The details of this structure, as well as how we generate our instance parameters in our study are given in A. The goal is to find a network design that minimizes the construction cost and the expected distribution cost.
@@ 3.2 Mixed Integer Programming Formulation
We model uncertainty with finitely many scenarios $\omega\in\Omega\coloneqq\{1,2,\ldots,W\}$. Each scenario $\omega$ has a probability of occurrence of $p_{\omega}$. Under scenario $\omega$, $s_{it}^{k\omega}$ represents the net supply of node $i$ for commodity $k$ in period $t$. Let $x_{at}^{k\omega}$ be the flow amount of commodity $k$ on arc $a$ on period $t$ under scenario $\omega$, which costs $c_{at}^{k}$ per unit, which we set linearly proportional to the arc length. Let $I_{it}^{k\omega}$ be the inventory level of node $i$ for commodity $k$ at the beginning of period $t\in\mathscr{T}\cup\{T\}$ under scenario $\omega$, which cannot exceed the inventory capacity of $C_{it}^{k\omega}$. Consistent with $s_{it}^{k\omega}$, we treat $C_{it}^{k\omega}$ as a scenario-dependent parameter, reflecting exogenous uncertainty in future infrastructure deployment. Capacity levels depend on the evolving role of each location in the hydrogen economy; for example, future suppliers are more likely to be assigned larger storage units. Holding costs, which typically include depreciation, risk of damage, and opportunity costs (Fallah, 2011), are less accounted for in hydrogen applications (HEAVENN, 2022). Thus, we assume no inventory holding costs in this study. However, the model can be adjusted to include these costs if necessary, without affecting the methodology presented in this paper. Let $e_{it}^{k\omega}$ be the amount of unsatisfied net demand in node $i$ for commodity $k$ in period $t$ under scenario $\omega$, which is penalized with a cost of $h$ per unit. This penalty may, for example, represent the variable cost of emergency transport by truck. Let the sets $\delta^{+}(i)$ and $\delta^{-}(i)$ represent the arc sets that enter and leave node $i$, respectively.
@@
The objective function minimizes the total cost of arc construction, arc conversion, the expected cost of arc flow, and the expected penalty cost. Constraint (2) ensures the inventory levels are balanced between periods according to the net supply, as well as the inflow and outflow of the commodity. Constraint (3) ensures that the inventory respects the capacity. Constraint (4) ensures that the flow capacity of an arc is not exceeded, and is only available for the assigned commodity. Constraint (5) ensures each arc can be built once in the planning horizon. Constraint (6) ensures that at the period an arc is constructed, it is assigned to its corresponding commodity. In the following periods, the arc may be converted for the use of another commodity. For those periods, Constraint (7) satisfies that each constructed arc is assigned to one of the commodities. We link the variable $z$ with the conversion variable $b$ by Constraint (8), ensuring the conversion cost is reflected in the objective function value. Each arc is restricted to be converted at most once during the planning horizon by Constraint (9). The initial networks for each commodity are created by Constraint (10). Finally, with Constraint (11) we make sure that the initial inventories are assigned to the start inventory values of $o_{i}^{k}$. The remaining constraints model the decision variables’ domains. An overview of the notation is given in Table 1.
Table 1: Overview of notation
Parameters:
$f_{at}^{k}$ fixed construction cost of pipeline $a$ at period $t$ for commodity $k$
$g_{at}^{k}$ fixed conversion cost of pipeline $a$ at period $t$ for commodity $k$
$p_{\omega}$ probability of occurrence of scenario $\omega$
$c_{at}^{k}$ unit flow cost of pipeline $a$ at period $t$ for commodity $k$
$h$ unit penalty cost of unsatisfied demand
$s_{it}^{k\omega}$ net supply of node $i$ at period $t$ for commodity $k$ under scenario $\omega$
$C_{it}^{k\omega}$ inventory capacity of node $i$ at period $t$ for commodity $k$ under scenario $\omega$
$U_{a}$ fixed flow capacity of pipeline $a$
Decision variables:
$y_{at}^{k}$ binary decision variable equaling $1$ if pipeline $a$ is constructed at period $t$ for commodity $k$
$b_{at}^{k}$ binary decision variable equaling $1$ if pipeline $a$ is converted at period $t$ for commodity $k$
$z_{at}^{k}$ binary decision variable equaling $1$ if pipeline $a$ is assigned at period $t$ for commodity $k$
$x_{at}^{k\omega}$ flow amount of pipeline $a$ at period $t$ for commodity $k$ under scenario $\omega$
$e_{it}^{k\omega}$ unsatisfied demand of node $i$ at period $t$ for commodity $k$ under scenario $\omega$
$I_{it}^{k\omega}$ inventory level of node $i$ at the beginning of period $t$ for commodity $k$ under scenario $\omega$
@@ 4 Progressive Hedging Based Matheuristic
$b_{at}^{k}\in\{0,1\}$ $\forall a\in\mathscr{A},k\in\mathscr{K},\omega\in\Omega,t\in\mathscr{T},$ (40)
As proposed by Rockafellar and Wets (1991), we relax the non-anticipativity constraints by removing Constraints (35) - (40), allowing the model for each $\omega\in\Omega$ to be solved individually. This removes the requirement for first-stage decisions to equal a single global variable (e.g., $y_{at}^{k}$). Instead, the Progressive Hedging algorithm encourages consensus by penalizing each scenario’s decisions for deviating from their weighted average. These averages are given below:
$\overline{y}_{at}^{k}\coloneqq\sum_{\omega\in\Omega}p_{\omega}y_{at}^{k\omega}$ $\forall a\in\mathscr{A},k\in\mathscr{K},t\in\mathscr{T},$ (41)
$\overline{b}_{at}^{k}\coloneqq\sum_{\omega\in\Omega}p_{\omega}b_{at}^{k\omega}$ $\forall a\in\mathscr{A},k\in\mathscr{K},t\in\mathscr{T},$ (42)
$\overline{z}_{at}^{k}\coloneqq\sum_{\omega\in\Omega}p_{\omega}z_{at}^{k\omega}$ $\forall a\in\mathscr{A},k\in\mathscr{K},t\in\mathscr{T}.$ (43)
Following the framework of Rockafellar and Wets (1991), The objective for each scenario $\omega$ is then modified with the following Lagrangian coefficients:
@@
Here, $\lambda$, $\delta$, and $\pi$ are Lagrangian multipliers for decision variables $y$, $b$, and $z$ respectively. The ratio $\rho$ is used to enforce non-anticipativity by penalizing the deviations of decision variables from their average values across scenarios, as well as acting as a step size for the updates on multipliers. Since all the first-stage variables are binary in our model, the quadratic objective function (44) would be equal to the following linearized objective function for all binary solutions:
@@
Algorithm 1 Progressive Hedging
@@
We apply four accelerating techniques to the base PHA in our solution method. The first acceleration technique is to consider bundles of scenarios instead of single-scenario subproblems. This results in solving two-stage recourse models (i.e., subject to the scenarios in the bundle) in the PHA, instead of solving deterministic single scenario problems. It has been shown that for well-calibrated bundle sizes, this reduces computation time without affecting the quality too much (Crainic et al., 2014). For this study, we use random assignment to form scenario bundles, serving as a baseline strategy to test the effectiveness of bundling. More sophisticated bundling strategies are left for future work, see in Section 7. We use a set of bundles $\Xi$ where each scenario $\omega\in\Omega$ is in exactly one bundle $\xi\in\Xi$. The probability of a bundle is assigned as $p_{\xi}=\sum_{\omega\in\xi}p_{\omega}$. The objective function of the two-stage recourse model for bundle $\xi$ is given below in Equation (46).
@@
We use a similar acceleration technique with two main differences. Firstly, we enable fixing variables even for the cases where they are not equal for all scenarios but for the majority of scenarios, i.e., with a soft fixing scheme. If the total probability of scenarios where the decision variable shares a common value exceeds an input probability parameter $p_{\textsc{h}}$, we fix the corresponding decision variables. Secondly, we observe that even if some decision variables do not share the same value, we may see similarities between the values of decision variables in terms of our problem context. Thus, we implement a problem-specific convergence rule that does not depend on individual decision variables as in Watson and Woodruff (2011). Instead, the convergence condition depends on subsets of decision variables. For example, consider a case where one arc is built for the majority of scenarios from node $i_{1}$ to $i_{2}$ for commodity $k$. However, since we allow multiple arcs between a node pair $(i_{1},i_{2})$ —each representing a pipeline with a different capacity— arcs with different capacities, or in different periods, may be built in different scenarios. In this case, the individual values of $y_{at}^{k\omega}$ suggest the algorithm is far from being converged, but the following would be satisfied for the majority of scenarios:
@@
To implement this logic of convergence, we create the acceleration technique given in Algorithm 2. We refer to this strategy as aggregate soft fixing, reflecting both its use of aggregated variable sums and its soft enforcement based on a probability threshold, $p_{\textsc{h}}$. The algorithm checks for Constraint (47) which is defined for decision variable $y$. We observe in preliminary experiments that the complexity of our model is mainly driven by the $y$ variables. As such, there is no need to consider the same logic for the variables $z$ and $b$. Indeed, they can be found rather easily for a fixed set of $y$ variables. This procedure is run once at each iteration of PHA.
This approach offers three main advantages. First, it enables earlier fixing than hard fixing techniques, since hard fixing requires $100\%$ agreement across scenarios; in contrast, aggregate soft fixing activates when the total probability of agreeing scenarios exceeds a threshold. Second, by focusing on aggregated subsets rather than individual variables, convergence is detected earlier and more broadly. Third, this aggregation leads to looser constraints—such as $\sum y\geq 1$ instead of strict equalities like y=1y=1—giving the solver more flexibility and helping it maintain near-optimality. Finally, while tailored to our network design context, this idea is not problem-specific: many two-stage recourse models in operations research contain similar aggregate patterns. We believe this makes aggregate soft fixing a promising extension for other progressive hedging applications, and we highlight this in Section 7 as a direction for future work.
Algorithm 2 softFix
@@
The third acceleration technique enables early termination of the PHA. As mentioned above, the majority of first-stage decision variables converge rather fast in the PHA. Often, the algorithm spends the majority of the iterations for some small discrepancies in the last decision variables. In order to accelerate, we terminate the algorithm before it is fully converged, and solve the rest of the problem with a MIP solver. One way of doing this would be fixing a subset of first-stage decision variables, and then solving the MIP for the rest. However, we implement a similar idea to the aggregate soft fixing and create heuristic constraints without fixing individual decision variables. It has the same heuristic constraint structure of Algorithm 2, but for early convergence, we also check the cases where Constraint (47) sums to zero. At each iteration, Algorithm 3 is executed in our method. Once the number of these heuristic constraints reaches the target stopping condition, we solve a MIP for the base model including these constraints.